OpenStudy (anonymous):

Please help: Solve the following system of equations by graphing. x – y = 7 x + y = 5 What is the x-value of their intersection? 6 -1 12 none of the above

OpenStudy (campbell_st):

add the 2 equations and y will be eliminated x - y = 7 x + y = 5 ---------- 2x = 12 divide by 2 to find x when you have x, substitute it into either equation to find y hope this helps.

OpenStudy (anonymous):

This helped me so much! thank you

OpenStudy (anonymous):

so x= 6. and then y would be x+y=5 6+y=5 -6 -6 y=-1 is that right?

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

how do i do that on x + 3y = 4 2x + 6y = 8 ? i thought you would multiply the top line by -2, but don't you have to do that on both lines?

OpenStudy (campbell_st):

thats the right method but they are the same line.... so they have an infinite number of common points this question has been written to see if you can recognize that fact.

OpenStudy (anonymous):

O okay. I get it. Thank you